OpenStudy (anonymous):

-5w+3/2=-6/5w-1/2 what is w

OpenStudy (anonymous):

\[\color{blue}{\text{Is it like this..}-5w+\frac{3}{2}=-\frac{6}{5w}-\frac{1}{2}}\]